### v1

@@ Abstract
Intelligence lies not only in the brain but in the body. The shape of our bodies can influence how we think and interact with the physical world. In robotics research, interacting with the physical world is crucial as it allows robots to manipulate objects in various real-life scenarios. Conventional robotic manipulation strategies mainly rely on finger-shaped end effectors. However, achieving stable grasps on fragile, deformable, irregularly shaped, or slippery objects is challenging due to difficulties in establishing stable force or geometric constraints.
Here, we present surface-based manipulation strategies that diverge from classical grasping approaches, using with flat surfaces as minimalist end-effectors. By changing the position and orientation of these surfaces, objects can be translated, rotated and even flipped across the surface using closed-loop control strategies. Since this method does not rely on stable grasp, it can adapt to objects of various shapes, sizes, and stiffness levels, even enabling the manipulation the shape of deformable objects. Our results provide a new perspective for solving complex manipulation problems.
@@ 1 Introduction
Physical morphologies of a body significantly influence how it interacts with and perceives its environment [1, 2, 3, 4, 5, 6]. In nature, many animals have evolved diverse grasping mechanisms to adapt to various contexts—for example, manual grippers resembling anthropomorphic hands [7], spinal grippers such as the prehensile tails of monkeys and certain lizards, and muscular hydrostats like octopus tentacles and elephant trunks [8]. Roboticists draw inspiration from the biological systems to develop grippers that utilize mechanical interlocking for geometric constraints [9, 10, 11], friction for force constraints [12, 13, 14, 15], or adhesion for stabilization [16, 17, 18]. These methods form the basis of most current robotic gripper designs [19, 20]. However, they often face challenges when manipulating irregularly shaped, deformable, or fragile objects [21]. Purely geometric constraints struggle with irregular shapes, and force constraints can damage deformable or delicate items. Reflecting on our daily experiences (see Fig. 1a), we often address similar challenges by adopting surface-based, nonprehensile manipulation methods and providing support from below. For example, while chopsticks are effective for picking up solid foods like fried potatoes, a spoon is more suitable for deformable substances like mashed potatoes. We hold a basketball with our palms rather than gripping it solely with our fingers, and we support a slippery fish from underneath. These examples demonstrate that when traditional grasping is ineffective, we naturally resort to surface-based, nonprehensile manipulation strategies. Motivated by this observation, introducing surface-based end effectors into robotics by adapting their morphology to planar forms offers a promising solution to these challenges.
@@
Figure 1: Surface-based manipulation. (a) In object manipulation contexts, grasping-based interactions such as using chopsticks or hand-grabbing a baseball form our intuition. However, surface-based manipulation plays an equally vital role. For instance, mashed potatoes are typically consumed with a spoon, a surface tool. Similarly, basketballs are often held or supported with palms. Moreover, using hands as a surface support from below makes it easier to hold a wet, slippery fish. For objects with non-conforming shapes or slippery surfaces, utilizing a nonprehensile, surface-based manipulation emerges as an intuitive choice. Building upon this, we present a novel approach to object manipulation using surfaces. This is achieved through the integration of three main motion principles: (b) translation, (c) flipping, and (d) rotation. (e) An example is in food packaging automation, where it addresses the challenge of handling items with varied shapes without causing damage. Surface-based manipulation enables tasks like rotating or positioning grape bunches for packaging and inspection.
@@
Robotics research has introduced devices capable of dynamically altering their surface geometries through actuation mechanisms [22, 23, 24, 25, 26, 27, 28, 29]. By adjusting their configurations, these systems can adapt to and interact more effectively with various objects. One approach employs grids of individually actuated parallel robots or linear actuators to create distributed manipulation surfaces [30, 31, 24, 32]. Another utilizes origami-inspired structures integrated into robotic gripper surfaces to modulate contact friction [33]. To enhance the capabilities of reconfigurable surfaces, modular designs incorporating soft actuators powered by vacuum [34] or origami-inspired actuators [35] have been explored. An alternative approach to address complex manipulation tasks involves integrating nonprehensile manipulation strategies that impose unilateral constraints on the object [36] and manipulating it with proper motion commands mainly based on machine learning principles [37]. A similar methodology includes two serial manipulators to perform complex tasks such as stir-frying which requires a sequence of synchronized and rapid movements [38]. Hu et al. used pneumatic morphological transformation to adjust wettability and manipulate droplets on a surface undergoing deformation [39]. Surface-based manipulation techniques have also been applied at smaller scales, specifically in the reconfigurable braille displays. These displays use dielectric elastomers [40], electromagnetism [41], or shape-memory alloys [42]. Furthermore, reconfigurable surfaces find applications in shape-changing displays designed for human interaction [23, 26, 25, 29].
@@
Despite these advancements, the practical application of reconfigurable surfaces in complex manipulation tasks remains limited. While the potential has been demonstrated in manipulating objects with simple geometries [43, 44, 45], challenges such as limited workspace and slow deformation speeds pose an obstacle for broader scenarios. Additionally, the necessity for continuous surface deformation reduces force output and limits their utility in manipulation contexts. These strategies frequently require the coordinated operation of many modules to manipulate a single object, complicating scalability to larger module sizes and quantities. Consequently, current applications lack dexterous strategies for handling non-rigid or irregularly shaped items and are often confined to specific hardware platforms. These limitations emphasize the need for novel design and actuation methodologies. Exploring innovative approaches to reconfigurable surfaces could significantly advance robotic manipulation, expanding robots’ capabilities in handling a diverse range of objects.
@@
Current nonprehensile manipulation methods on reconfigurable surfaces, which utilize individually actuated units, rely heavily on the kinematics of the modules and the sizes of the objects being manipulated. Together, these factors determine the types of nonprehensile motion primitives that can be utilized. Consequently, developing a comprehensive framework to analyze and compare these platforms is crucial for optimizing their design and manipulation strategies. Using information obtained from published literature on several platforms [46, 34, 24, 32, 47, 48, 31, 49, 50], we developed a scaling model for surface-based manipulation that results in the mapping of mobility and manipulability of these platforms. Finding the optimal setup for surface-based manipulation involves determining the exact number of DoF needed for each module and how many modules are necessary for different manipulation tasks. Table 1 lists various surface-based object manipulation examples and summarizes their capabilities in terms of repositioning and reorienting objects placed on them. We standardize the module dimensions across the platforms to facilitate comparative analysis, study the relationship between the working modules and manipulability, and highlight their capability in accommodating objects of diverse sizes. The minimum manipulation workspace, $MW_{min}$, indicates the smallest area where both the functions of the reconfigurable surface and the manipulation strategies are effectively employed. Then, we calculate the minimum mobility requirement, $MR_{min}$, that represents the necessary number of actuators for different types of surface-based manipulation techniques using single unit DoF, $DoF_{SU}$, and surface size, $SS_{SU}$:
@@
WaveHandling [46] 1 $d\times d$ >$d\times d$ $2d\times 2d$ 4
Soft Surface [34] 1 $d\times d$ >$2d\times 2d$ $3d\times 3d$ 9
inFORM [24] 1 $d\times d$ >$2d\times 2d$ $3d\times 3d$ 9
ArrayBot [32] 1 $d\times d$ >$2d\times 2d$ $3d\times 3d$ 9
Omnia Wheel [47] 3 $d\times d$ >$2d\times 2d$ $3d\times d$ 9
Planar Manip. [48] 3 $d\times d$ <$d\times d$ $d\times d$ 3
Delta Arrays [31] 3 $d\times d$ >$2d\times 2d$ $3d\times 3d$ 27
Soft Table [49] 4 $d\times d$ >$2d\times 2d$ $3d\times 3d$ 36
RoDyMan [50] 12 $d\times d$ <$d\times d$ $d\times d$ 12
@@
Surfaces consisting of multiple units with one $DoF_{SU}$ can form different profiles from a flat surface, such as slopes for rolling objects or ridges for flipping them, and they enable manipulation of objects larger than $SS_{SU}$ [24, 34, 32, 46]. In these examples, the $MR_{min}$ and consequently the actuator number typically remain relatively high, up to 900, mainly due to the necessary motion primitives and the properties of the manipulated objects [24, 34, 32, 46]. Increasing the $DoF_{SU}$ enables the utilization of additional motion primitives with fewer units. For instance, an example system comprises three $DoF_{SU}$ arranged in an 8×8 configuration [31, 30]. These units, having finger-like end effectors, manipulate objects through coordinated movements such as sliding in addition to rolling [31, 30]. Similarly, Deng et al. introduced a pneumatically actuated soft modular surface with four $DoF_{SU}$ capable of translating and rotating objects through small deformations on the surface [49] and complementary trajectory planning [51]. Although the number of units is lower in examples with higher $DoF_{SU}$, the $MR_{min}$ remains high due to the increased actuator number requirements. The introduction of dynamic movements such as vibration has also been shown to facilitate both translation and rotation within the plane [48]. Utilizing a similar methodology, surface-based manipulation extends to serial manipulators through the use of flat-surface end effectors. Ruggiero et al. utilized a dual-arm robot with a total of 12 DoF and employed dynamic manipulation techniques such as rolling, tossing, and batting that translate or rotate objects [36]. Analyzing the $MR_{min}$ aspects of the state-of-the-art reveals a trade-off between the number of required modules and the $DoF_{SU}$. For instance, a higher $DoF_{SU}$ for fewer units to accomplish the same task. However, increasing the $DoF_{SU}$ poses challenges due to hardware limitations, making it more difficult to expand the $SS_{SU}$ by adding more modules. Consequently, we adopt two units with three $DoF_{SU}$ that represent the minimum configuration required to construct a reconfigurable surface capable of performing dexterous manipulation tasks. This surface consists of individual elements arranged in flat configuration that is able to form three-dimensional configurations and provide synchronized movements upon actuation. The three $DoF_{SU}$, along with the utilization of the interaction between the two modules, enable a range of dynamic manipulation techniques. This configuration ensures the lowest $MR_{min}$ for surface-based manipulation platforms that can perform object translation, rotation, and flipping (see Fig. 1b-d). The system exhibits scalability in terms of module size, module number, and the size of the object being manipulated. Depending on the specific requirements of the objects, the proposed strategies can be adopted to accommodate different module sizes. Furthermore, increasing the number of modules allows for the application of these strategies to multiple objects simultaneously.
@@
We propose manipulation strategies based on reconfigurable modular surfaces to perform fundamental manipulation tasks such as translating, rotating, and flipping objects of various forms, including non-rigid and irregularly shaped items. We demonstrate these strategies using two identical modular and foldable robots placed side by side to form a reconfigurable surface operating through a decoupled actuation procedure. Each robot is designed, manufactured, and assembled using an origami-based approach combined with the Canfield joint principle [52, 35, 53, 54, 55] and has three DoF. The hierarchical closed-loop control architecture overcomes kinematic constraints and adapts to different robotic platforms, which addresses hardware adaptability issues. By eliminating the need for grasping and continuous surface deformation, our method enables the manipulation of objects with diverse shapes and forms. Quantitative closed-loop experiments validate the overall performance of our system and demonstrate that our approach achieves more dexterous tasks with fewer degrees of freedom compared to other surface-based methods. This work also shows the potential of combining different manipulation strategies to alter the shape of flexible and deformable objects, overcoming existing limitations and expanding the capabilities of robotic manipulation systems. Our approach offers a new perspective for applications such as food packaging (see Fig. 1e) involving reconfigurable surfaces, where automation faces the challenge of handling items with varied and unpredictable shapes without causing damage.

### v2

@@ Abstract
Intelligence lies not only in the brain (decision-making processes) but in the body (physical morphology). The morphology of robots can significantly influence how they interact with the physical world, crucial for manipulating objects in real-life scenarios. Conventional robotic manipulation strategies mainly rely on finger-shaped end effectors. However, achieving stable grasps on fragile, deformable, irregularly shaped, or slippery objects is challenging due to difficulty in establishing stable forces or geometric constraints.
Here, we present surface-based manipulation strategies that diverge from classical grasping approaches, using flat surfaces as minimalist end-effectors. By adjusting surfaces’ position and orientation, objects can be translated, rotated, and flipped across the surface using closed-loop control strategies. Since this method does not rely on stable grasping, it can adapt to objects of various shapes, sizes, and stiffness levels, even enabling manipulation of the shape of deformable objects. Our results provide a new perspective for solving complex manipulation problems.
@@ 1 Introduction
Physical morphologies significantly shape how a system interacts with and perceives its environment [embodiment, sun2023embedded, baines2024robots, shah2021soft, baines2022multi, gupta2021embodied]. Morphologies that are structurally adapted to a task reduce the complexity of control and enable more precise and responsive task execution. In nature, many animals have evolved diverse grasping mechanisms to adapt to various contexts—for example, manual grippers resembling anthropomorphic hands [pouydebat2023convergent], spinal grippers such as the prehensile tails of monkeys and certain lizards, and muscular hydrostats like octopus tentacles and elephant trunks [softgrip]. Roboticists draw inspiration from the biological systems to develop grippers that utilize mechanical interlocking for geometric constraints [katzschmann, hao_2023, howard_2022], friction for force constraints [7390277, 9000610, 10146043, 7994658], or adhesion for stabilization [Chen_2021, shintake_2016, piskarev_2023]. These methods form the basis of most current robotic gripper designs [ZHANG2020105694]. However, they often face challenges when manipulating irregularly shaped, deformable, or fragile objects [doi:10.1126/science.aat8414]. Purely geometric constraints struggle with irregular shapes, and force constraints can damage deformable or delicate items. To address these challenges, researchers have developed advanced gripper designs that enhance robotic manipulation. Inspired by biological systems, soft robotic grippers with mechanical compliance can handle fragile objects [SoftActuators1, softgriperreview]. Granular jamming grippers conform to objects of arbitrary shape to enable their pick-and-place operations [GranularJamming1, GranularJamming2], while tactile sensing improves real-time adjustments for delicate objects [tactilesensing, GranularJamming3]. These methods still depend on stable grasps that match the size of the object, and struggle with deformable, flat, or highly variable objects [softgriperreview]. Reflecting on our daily experiences (see Fig. 1a), we often address similar challenges by adopting surface-based, nonprehensile manipulation methods and providing support from below. For example, while chopsticks are effective for picking up solid foods like fried potatoes, a spoon is more suitable for deformable substances like mashed potatoes. We hold a basketball with our palms rather than gripping it solely with our fingers, and we support a slippery fish from underneath. These examples demonstrate that when traditional grasping is ineffective, we naturally resort to surface-based, nonprehensile manipulation strategies. Motivated by this observation, introducing surface-based end effectors into robotics by adapting their morphology to planar forms offers a promising solution to these challenges.
@@
Figure 1: Surface-based manipulation. (a) Object properties influence whether we intuitively grasp or support them with a surface. Small, rigid objects with defined edges are easily grasped, while large, round, deformable, soft, or slippery objects are more effectively manipulated through surface contact. This distinction highlights how surface-based manipulation complements traditional grasping in handling diverse objects. Building upon this, we present a novel approach to robotic object manipulation using surfaces. This is achieved through the integration of three main motion principles: (b) translation, (c) flipping, and (d) rotation. (e) An example is in food packaging automation, where it addresses the challenge of handling items with varied shapes without causing damage. Surface-based manipulation enables tasks like rotating or positioning grape bunches for packaging and inspection.
@@
Robotics research has introduced devices capable of dynamically altering their surface geometries through actuation mechanisms [rasmussen2012shape, alexander2018grand, follmer2013inform, gronbaek2020kirigamitable, steed2021mechatronic, hwang2022shape, tahouni2020nurbsforms, siu2018shapeshift]. By adjusting their configurations, these systems can adapt to and interact more effectively with various objects. One approach employs grids of individually actuated parallel robots or linear actuators to create distributed manipulation surfaces [Thompson2021, patil2023, follmer2013inform, xue2023arraybot]. Another utilizes origami-inspired structures integrated into robotic gripper surfaces to modulate contact friction [Lu2020c]. To enhance the capabilities of reconfigurable surfaces, modular designs incorporating soft actuators powered by vacuum [Robertson2018] or origami-inspired actuators [Salerno2020] have been explored. Hu et al. used pneumatic morphological transformation to adjust wettability and manipulate droplets on a surface undergoing deformation [hu2022]. Surface-based manipulation techniques have also been applied at smaller scales, specifically in the reconfigurable braille displays. These displays use dielectric elastomers [Qu2021], electromagnetism [Bettelani2020], or shape-memory alloys [Haga2005]. Furthermore, reconfigurable surfaces find applications in shape-changing displays designed for human interaction [alexander2018grand, steed2021mechatronic, gronbaek2020kirigamitable, siu2018shapeshift, stanley2016closed]. An alternative approach to address complex manipulation tasks involves integrating nonprehensile manipulation strategies that impose unilateral constraints on the object [ruggiero_2018]. For instance, [R2cite2] demonstrates a paddle-like end-effector with visual feedback for pick-and-place tasks on rigid hexahedral objects, while [R2cite3] employs two 6 degrees of freedom (DoF) robotic arms with dual “palms” for reorientation tasks. In [R2cite1, R2cite6], a single motor-driven flexible joint mechanism enables planar manipulation via vibration modes. Inspired by pizza paddles, [R2cite4, R2cite5] introduce planar end-effectors facilitating object translation and rotation, and a related method uses two serial manipulators for rapid, coordinated actions like stir-frying [liu_2022]. While some of these systems incorporate visual feedback [R2cite2, R2cite1] or explore deformable object manipulation [R2cite5, higashimori2013gait], many lack closed-loop control, focus on limited range of rigid objects, or rely on single-surface methods that limit control over contact forces and scalability.
@@
Modular reconfigurable surface-based strategies also face significant practical limitations. While the potential has been demonstrated in manipulating spheres [liu2021robotic, johnson2023multifunctional], challenges such as limited workspace and slow deformation speeds pose an obstacle for broader scenarios. Additionally, the necessity for continuous surface deformation reduces force output and limits their utility in manipulation contexts. These strategies frequently require the coordinated operation of many modules to manipulate a single object, complicating scalability to larger module sizes and quantities. Consequently, current applications lack dexterous strategies for handling non-rigid or irregularly shaped items and are often confined to specific hardware platforms. These limitations emphasize the need for novel design and actuation methodologies. Exploring innovative approaches to reconfigurable surfaces could significantly advance robotic manipulation, expanding robots’ capabilities in handling a diverse range of objects.
@@
Current nonprehensile manipulation methods on reconfigurable surfaces, which utilize individually actuated units, rely heavily on the kinematics of the modules and the sizes of the objects being manipulated. Together, these factors determine the types of nonprehensile motion primitives that can be utilized. Consequently, developing a comprehensive framework to analyze and compare these platforms is crucial for optimizing their design and manipulation strategies. Using information obtained from published literature on several platforms [wavehandling, Robertson2018, follmer2013inform, xue2023arraybot, R2cite2, R2cite3, R2cite1, R2cite4, omniawheel, platemanipulator, patil2023, softmachinetable, rodyman], we developed a scaling model for surface-based manipulation that results in the mapping of mobility and manipulability of these platforms. Finding the optimal setup for surface-based manipulation involves determining the exact number of DoF needed for each module and how many modules are necessary for different manipulation tasks. Table 1 lists various surface-based object manipulation examples and summarizes their capabilities in terms of repositioning and reorienting objects placed on them. We standardize the module dimensions across the platforms to facilitate comparative analysis, study the relationship between the working modules and manipulability, and highlight their capability in accommodating objects of diverse sizes. The minimum manipulation workspace, $MW_{min}$, indicates the smallest area where both the functions of the reconfigurable surface and the manipulation strategies are effectively employed. Then, we calculate the minimum mobility requirement, $MR_{min}$, that represents the necessary number of actuators for different types of surface-based manipulation techniques using single unit DoF, $DoF_{SU}$, and surface size, $SS_{SU}$:
@@
Single Act. [R2cite1] 1 $\times d <d×d$$\times d$ $d\times d$ 1
WaveHandling [wavehandling] 1 $\times d >d×d$$\times d$ $2d\times 2d$ 4
Soft Surface [Robertson2018] 1 $d\times d$ $2d\times 2d$ $3d\times 3d$ 9
inFORM [follmer2013inform] 1 $d\times d$ $2d\times 2d$ $3d\times 3d$ 9
ArrayBot [xue2023arraybot] 1 $d\times d$ $2d\times 2d$ $3d\times 3d$ 9
Pizza Peel [R2cite4] 2 $\times d <d×d$$\times d$ $d\times d$ 2
Omnia Wheel [omniawheel] 3 $d\times d$ $2d\times 2d$ $3d\times d$ 9
Planar Manip. [platemanipulator] 3 $\times d <d×d$$\times d$ $d\times d$ 3
Delta Arrays [patil2023] 3 $d\times d$ $2d\times 2d$ $3d\times 3d$ 27
Soft Table [softmachinetable] 4 $d\times d$ $2d\times 2d$ $3d\times 3d$ 36
Dynamic Manip. [R2cite2] 6 $\times d <d×d$$\times d$ $d\times d$ 6
[…]
RoDyMan [rodyman] 12 $\times d <d×d$$\times d$ $d\times d$ 12
@@
Surfaces consisting of multiple units with one $DoF_{SU}$ can form different profiles from a flat surface, such as slopes for rolling objects or ridges for flipping them, and they enable manipulation of objects larger than $SS_{SU}$ [follmer2013inform, Robertson2018, xue2023arraybot, wavehandling]. In these examples, the $MR_{min}$ and consequently the actuator number typically remain relatively high, up to 900, mainly due to the necessary motion primitives and the properties of the manipulated objects [follmer2013inform, Robertson2018, xue2023arraybot, wavehandling]. Increasing the $DoF_{SU}$ enables the utilization of additional motion primitives with fewer units. For instance, an example system comprises three $DoF_{SU}$ arranged in an 8×8 configuration [patil2023, Thompson2021]. These units, having finger-like end effectors, manipulate objects through coordinated movements such as sliding in addition to rolling [patil2023, Thompson2021]. Similarly, Deng et al. introduced a pneumatically actuated soft modular surface with four $DoF_{SU}$ capable of translating and rotating objects through small deformations on the surface [softmachinetable] and complementary trajectory planning [softtable_traj]. Although the number of units is lower in examples with higher $DoF_{SU}$, the $MR_{min}$ remains high due to the increased actuator number requirements. The introduction of dynamic movements such as vibration has also been shown to facilitate both translation and rotation within the plane [platemanipulator]. Utilizing a similar methodology, surface-based manipulation extends to serial manipulators through the use of flat-surface end effectors. Ruggiero et al. utilized a dual-arm robot with a total of 12 DoF and employed dynamic manipulation techniques such as rolling, tossing, and batting that translate or rotate objects [ruggiero_2018]. Analyzing the $MR_{min}$ aspects of the state-of-the-art reveals a trade-off between the number of required modules and the $DoF_{SU}$. For instance, a higher $DoF_{SU}$ for fewer units to accomplish the same task. However, increasing the $DoF_{SU}$ poses challenges due to hardware limitations, making it more difficult to expand the $SS_{SU}$ by adding more modules. Consequently, we adopt two units with three $DoF_{SU}$ that represent the minimum configuration required to construct a reconfigurable surface capable of performing dexterous manipulation tasks. This surface consists of individual elements arranged in flat configuration that is able to form three-dimensional configurations and provide synchronized movements upon actuation. The three $DoF_{SU}$, along with the utilization of the interaction between the two modules, enable a range of dynamic manipulation techniques. This configuration ensures the lowest $MR_{min}$ for surface-based manipulation platforms that can perform object translation, rotation, and flipping (see Fig. 1b–d). While some prior work [R2cite2] achieves a similar $MR_{min}$, using a single platform with six DoF to perform comparable manipulation modes, the modularity in our system distributes the required MR across two smaller units, each with three DoF. This bimanual design enhances dexterity by enabling the simultaneous application of forces from multiple directions, allowing for more dexterous tasks such as folding deformable objects. It also improves scalability by enabling incremental expansion of functionality through additional modules, rather than increasing complexity within a single unit. Moreover, our system supports closed-loop control and quantitative evaluation and handles diverse forms of objects. The system exhibits scalability in terms of module size, module number, and the size of the object being manipulated. Depending on the specific requirements of the objects, the proposed strategies could be adopted to accommodate different module sizes. Furthermore, increasing the number of modules allows for the application of these strategies to multiple objects simultaneously.
@@
We propose manipulation strategies based on reconfigurable modular surfaces to perform fundamental manipulation tasks such as translating, rotating, and flipping objects of various forms, including non-rigid and irregularly shaped items. We demonstrate these strategies using two identical modular and foldable robots placed side by side to form a reconfigurable surface operating through a decoupled actuation procedure. Each robot is designed, manufactured, and assembled using an origami-based approach combined with the Canfield joint principle [Canfield_1998, Salerno2020, mintchev2019portable, Mete2021, robertson2021soft] and has three DoF. The hierarchical closed-loop control architecture overcomes kinematic constraints and adapts to different robotic platforms, which addresses hardware adaptability issues. By eliminating the need for grasping and continuous surface deformation, our method enables the manipulation of objects with diverse shapes and forms. Quantitative closed-loop experiments validate the overall performance of our system and demonstrate that our approach achieves more dexterous tasks with fewer DoF compared to other surface-based methods. This work also shows the potential of combining different manipulation strategies to alter the shape of flexible and deformable objects, overcoming existing limitations and expanding the capabilities of robotic manipulation systems. Our approach offers a new perspective for applications such as food packaging (see Fig. 1e) involving reconfigurable surfaces, where automation faces the challenge of handling items with varied and unpredictable shapes without causing damage.